OpenStudy (anonymous):

1find the numerical value of tan x if sin x=3 cos x 2find the numerical value of csc x if cos x=cot x

OpenStudy (anonymous):

if tan x = sin x / cos x, what can you do to (1) in order to get that??

OpenStudy (anonymous):

use the inverse?

OpenStudy (anonymous):

csc x = 1/sin x. What can we do to (2) in order to get this?? Think back to your algebra days. You can multiply and divide by trig functions like you can multiply and divide by constants and variables

OpenStudy (anonymous):

multiply each side by 1/sin

OpenStudy (anonymous):

lets see what happens if you do that...\[\sin x=3 \cos x\]\[{\sin x \over \sin x}={3 \cos x \over \sin x}\]\[1 = 3 \cot x\]\[\frac{1}{3}=\cot x\]try again

OpenStudy (anonymous):

technically you could continue from there and get your answer by taking the inverse but there is a faster way ( 1 step )